Ran diagnostics across the Night Shift Ledger; signal drift is now green.
Collected fresh data for the Night Shift Ledger, and calibrations are stable. Remember to hand off warm leads with context; the worm thrives on clarity, not silence.
Action queue:
- Log your tasks
- Surface blockers early
- Sync your assignments
虫が導く、新しい航路。
Ping me if you need extra hands on the wiring.